OpenStudy (anonymous):

I've learned this before .. i just need a recap or review to make sure i'm remembering it right. Find the first six terms of the sequence. a1 = 6, an = 2 • an-1

OpenStudy (anonymous):

a1 = 6, an = 2 • an-1 6, 12, 14, 16, 18, 20 12, 24, 48, 96, 192, 384 0, 2, 12, 14, 16, 18 6, 12, 24, 48, 96, 192

OpenStudy (anonymous):

Well: \[ \begin{align} a_1 &= 6\\ a_n &= 2a_{n-1} \end{align} \] So: \[ \begin{align} a_2 &= 2a_1 = 2 \cdot 6 = 12\\ a_3 &= 2a_2 = 2 \cdot 12 = 24 \end{align} \] And so on.

OpenStudy (anonymous):

Yeah alright see i got up to the second part and was doing 2*6 ... but now that you said that i remember learning it that way .. thank you so much .. mind helping with another ?

OpenStudy (anonymous):

Sure. Fire away.

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ... an = -17 + 5(n + 2) an = -17 + 5(n + 1) an = -17 + 5(n - 1) an = -17 x 5(n - 1)

OpenStudy (anonymous):

i think its B ?

OpenStudy (anonymous):

i could be completely wrong .. i know for sure its not D

OpenStudy (anonymous):

Either B or C ?

OpenStudy (anonymous):

As far as I can tell, neither. I think it should be: \[ a_n = -17 + 5n \]

OpenStudy (anonymous):

hmm thats strange .. ill go with b just because thats what popped out at me .. mind helping with 1 more ? sorry lol

OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -4, an = an-1 + 7 -4, 3, 10, 17, 24, 31 -4, 7, 14, 21, 28, 35 3, 10, 17, 24, 31, 38 0, 7, 14, 21, 28, 35 I think its A ?

OpenStudy (anonymous):

It has to be A or B, and since the second element should be 3, it's A. Remember: it's just algebra. a1 is a variable just like x is, and a2 is just something plus a1. a3 is just something plus a2, and so on.